# MonarchAttention: Zero-Shot Conversion to Fast, Hardware-Aware Structured Attention

### v1

@@ Abstract
Transformers have achieved state-of-the-art performance across various tasks, but suffer from a notable quadratic complexity in sequence length due to the attention mechanism. In this work, we propose MonarchAttention – a novel approach to sub-quadratic attention approximation via Monarch matrices, an expressive class of structured matrices. Based on the variational form of softmax, we describe an efficient optimization-based algorithm to compute an approximate projection of softmax attention onto the class of Monarch matrices with $\Theta(N\sqrt{N}d)$ computational complexity and $\Theta(Nd)$ memory/IO complexity. Unlike previous approaches, MonarchAttention is both (1) transferable, yielding minimal performance loss with no additional training, even when replacing every attention layer of the transformer, and (2) hardware-efficient, utilizing the highest-throughput tensor core units on modern GPUs. With optimized kernels, MonarchAttention achieves substantial speed-ups in wall-time over FlashAttention-2: 1.4×1.4$\times$1.4 × for shorter sequences $(N=256)$, 4.5×4.5$\times$4.5 × for medium-length sequences $(N=4K)$, and 8.2×8.2$\times$8.2 × for longer sequences $(N=16K)$. We demonstrate the quality of MonarchAttention on diverse tasks and architectures in vision and language problems, showing that it flexibly and accurately approximates softmax attention in a variety of contexts. Our code is available at https://github.com/cjyaras/monarch-attention.
@@ 1 Introduction
Over the past decade, transformers (Vaswani et al., 2017) have become the dominant architecture for generating and processing various data modalities, such as text (Brown et al., 2020), images (Dosovitskiy et al., 2021), and speech (Radford et al., 2023). Central to the transformer’s success is attention, the mechanism through which complex interactions within sequential data are captured through weighted combinations of embeddings at every position in the sequence. Famously, the attention mechanism has a quadratic-time complexity $\Theta(N^{2}d)$ in the length of the sequence $N$, where $d$ is the head dimension, which is a key bottleneck for both training and inference, particularly in long sequence problems. To address this, numerous works have proposed sub-quadratic substitutes for attention. Yet, such approaches either (1) are not transferable, requiring training from scratch or fine-tuning of existing models, or (2) do not yield speed-ups in practice (except on extremely long sequences) due to a gap between theoretical complexity and practical considerations for modern GPUs, especially compared to highly optimized implementations (Dao et al., 2022b).
In this work, we propose MonarchAttention: a novel sub-quadratic attention substitute based on approximating the attention matrix via Monarch matrices (Dao et al., 2022a), a class of expressive structured matrices. At first glance, this is computationally infeasible – for sequence length $N$, computing an exact projection onto the set of Monarch matrices has a super-quadratic $O(N^{2}\sqrt{N})$-time complexity, not to mention that we need to form the entire $N\times N$ attention matrix. Instead, we reframe the computation of the attention matrix as an optimization problem in terms of the variational form of softmax, and exploit low-dimensional structure in the variational objective when constrained to the set of Monarch matrices – this yields a sub-quadratic $\Theta(N\sqrt{N}d)$-time approximation, where $d$ is the head dimension. This approach is analogous to optimization-based approaches for low-rank approximation of a matrix (Chi et al., 2019), where rather than computing a full SVD and truncating to the desired rank, one can more efficiently minimize a Frobenius norm objective constrained to the set of low-rank matrices. We briefly review prior work on structured matrices, including Monarch matrices, as well as existing approaches to efficient attention.
@@
We note that there are significant drawbacks to the approaches described above. Pure low-rank methods are often fast and hardware-friendly, but are not typically suitable as drop-in replacements for attention in pre-trained transformers due to the prevalence of “strongly diagonal”, high-rank attention matrices where attention weights are concentrated locally in a sequence. Making up for this with a fixed sparsity pattern does not allow for data-dependent support of the attention matrix, necessary for zero-shot conversion. Finally, sparsity/LSH-based approaches that do not have a fixed sparsity pattern, improve on accuracy over low-rank approximations but suffer from significant overhead due to GPU incompatibility. MonarchAttention, on the other hand, achieves the best of both worlds: it is fast and hardware-friendly due to utilization of tensor cores for batched matmuls, while computing highly accurate approximations to the extent that it can directly replace softmax attention with no additional training.
@@ 2 Preliminaries
We use $[N]$ to denote the index set $\{1,2,\dots,N\}$. We use $\Delta^{N}$ to denote the $(N-1)$ dimensional unit simplex, given by $\Delta^{N}=\{\bm{a}\in\mathbb{R}^{N}:\bm{a}\succeq 0,\langle\bm{1}_{N},\bm{a}\rangle=1\}$. We denote the $m\times m$ identity matrix by $\bm{I}_{m}$. We use the notation $\bm{A}_{ijk}$ to denote an element of a 3-way tensor, and $\bm{A}_{i,:,k}$ to denote a slice. We use $\delta_{kl}$ to denote the Kronecker delta that is $1$ if $k=l$ and otherwise 0.
@@
The softmax function $\mathbb{R}^{N}\rightarrow\Delta^{N}$ maps $N$ real numbers to the $(N-1)$-dimensional unit simplex, and is defined as
$[\operatorname{softmax}(\bm{z})]_{i}:=\frac{\exp(\bm{z}_{i})}{\sum_{j}\exp(\bm{z}_{j})},\;\forall i\in[N].$ (1)
@@ 3 MonarchAttention
where $\Delta^{N\times N}$ denotes a matrix whose rows lie on $\Delta^{N}$, and $H(\bm{A})=-\sum_{i,j}\bm{A}_{ij}\log\bm{A}_{ij}$. For a dense matrix $\bm{A}$, computing $f(\bm{A};\bm{Q},\bm{K})$ requires $\Theta(N^{2}d)$ operations, which is the same as computing $\sigma(\bm{Q}\bm{K}^{\top})$ directly. However, we are interested in the case where $\bm{A}$ is a Monarch matrix $\bm{M}=\bm{P}^{\top}\bm{B}$:
$f(\bm{P}^{\top}\bm{B};\bm{Q},\bm{K})=\langle\bm{P}^{\top}\bm{B},\bm{Q}\bm{K}^{\top}\rangle+H(\bm{P}^{\top}\bm{B})=\langle\bm{B},\widetilde{\bm{Q}}\bm{K}^{\top}\rangle+H(\bm{B})=\sum_{j,k}f(\bm{B}_{jk};\widetilde{\bm{Q}}_{j},\bm{K}_{k}),$
where $\widetilde{\bm{Q}}=\bm{P}\bm{Q}$, and $\widetilde{\bm{Q}}_{j}\in\mathbb{R}^{m\times d},\bm{K}_{k}\in\mathbb{R}^{b\times d}$ are the $j$th and $k$th block of rows of $\widetilde{\bm{Q}},\bm{K}$ respectively. Then, for each $j\in[b],k\in[m]$ we evaluate $f$ on the rank-one matrix $\bm{B}_{jk}=\bm{L}_{jk}\bm{R}_{kj}^{\top}$:
@@
$=\langle\widetilde{\bm{Q}}_{j}^{\top}\bm{L}_{jk},\bm{K}_{k}^{\top}\bm{R}_{kj}\rangle-\left(\bm{1}^{\top}\bm{R}_{kj}\right)\cdot H(\bm{L}_{jk})-\left(\bm{1}^{\top}\bm{L}_{jk}\right)\cdot H(\bm{R}_{kj}).$
@@
i.e., slices of $\bm{L},\bm{R}$ individually lying on the unit simplex is sufficient to enforce the constraint on $\bm{M}$. This is easily seen from (5) – obviously if $\bm{L}_{jkl},\bm{R}_{kji}\geq 0$, then $[\bm{M}_{lk}]_{ji}\geq 0$. Moreover, this also enforces the sum-to-one constraint, as rows of $\bm{M}$ sum as
@@
We now present the updates for $\bm{L},\bm{R}$. Initializing $\bm{L}_{jkl}^{(0)}=\delta_{kl}$ as block identity, we have
@@
where $\overline{\bm{Q}}_{jl},\overline{\bm{K}}_{ki}\in\mathbb{R}^{d}$ are the $(b\cdot(l-1)+j)$th and $(b\cdot(k-1)+i)$th row of $\bm{Q}$ and $\bm{K}$ respectively. The full derivation is provided in Section C.1. After $T$ steps, we obtain the final Monarch approximation $\bm{M}^{(T)}\approx\sigma(\bm{Q}\bm{K}^{\top})$ with factors $\bm{L}^{(T)}$ and $\bm{R}^{(T)}$, from which we output $\bm{M}^{(T)}\bm{V}$ using (4). A naïve implementation of the full algorithm is provided in Section C.2. We discuss in Section C.3 how padding can be incorporated into MonarchAttention for when $N$ is not divisible by $b$.
@@
which can be seen as a batched attention computation, meaning we can implement a FlashAttention-like kernel to reduce IO between HBM and on-chip SRAM. However, several aspects of this computation make it particularly IO-efficient. Besides the fact that $\overline{\bm{K}}$ acts as both the $\bm{K}$ and $\bm{V}$ matrices in (3), the effective sequence length is $\sqrt{N}$. This eliminates the need for tiling along the sequence length, except for very long sequences having $\Theta(\sqrt{N}d)>S$, where $S$ is the size of on-chip SRAM. This means that we have an optimal IO complexity of $\Theta(Nd)$ for a single call, as opposed to the worst-case $O(N^{2}d^{2}/S)$ complexity of FlashAttention. The computation of $\bm{\alpha}_{R}^{(t)}$ from $\bm{\alpha}_{L}^{(t)}$, as well as the Monarch matmul, can be written in a similar fashion. Based on this, MonarchAttention not only achieves significant speed-up over FlashAttention for longer sequences, but also for shorter ones. Python-like code for MonarchAttention is given in Figure 4.
{minted}
[frame=single,linenos,fontsize=]python def al_cl_kernel(aR, cR, Kb): # Computes aL, cL from aR, cR R = softmax(bmm(aR, Kb.transpose(1, 2)) / cR[:, :, None], dim=2) cL = sum(R * log(R), dim=2).transpose(0, 1) aL = bmm(R, Kb).transpose(0, 1) return aL, cL
def ar_cr_kernel(aL, cL, Qb): # Computes aR, cR from aL, cL L = softmax(bmm(aL, Qb.transpose(1, 2)) - cL[:, :, None], dim=1) cR = sum(L, dim=2).transpose(0, 1) aR = bmm(L, Qb).transpose(0, 1) return aR, cR
def al_y_cl_kernel(aR, cR, Kb, Vb): # Fuse al_cl_kernel + Monarch matmul 1st step R = softmax(bmm(aR, Kb.transpose(1, 2)) / cR[:, :, None], dim=2) cL = sum(R * log(R), dim=2).transpose(0, 1) aL = bmm(R, Kb).transpose(0, 1) y = bmm(R, Vb).transpose(0, 1) return aL, y, cL
def z_kernel(aL, y, cL, Qb): # Monarch matmul 2nd step L = softmax(bmm(Qb, aL.transpose(1, 2)) - cL[:, None, :], dim=2) z = bmm(L, y).transpose(0, 1) return z
def monarch_attention(Q, K, V, T): # Q, K, V: (N, d), T: number of steps Qb = Q.reshape(m, b, d).transpose(0, 1) Kb = K.reshape(m, b, d) Vb = V.reshape(m, b, d) aR = Q.reshape(m, b, d) cR = ones(m, b)
for t in range(T-1): aL, cL = al_cl_kernel(aR, cR, Kb) aR, cR = ar_cr_kernel(aL, cL, Qb)
aL, y, cL = al_y_cl_kernel(aR, cR, Kb, Vb) z = z_kernel(aL, y, cL, Qb) o = z.reshape(N, d) return o
@@ 4 Experiments
In this section, we evaluate the zero-shot performance (no additional training) of MonarchAttention for converting pre-trained/fine-tuned transformer attention layers to sub-quadratic attention in four different model/task settings. We compare with previous low-rank attention methods (Katharopoulos et al., 2020; Choromanski et al., 2021; Xiong et al., 2021; Qin et al., 2022); see Section D.1 for more details on the baselines. We specifically exclude low-rank methods with learnable components (Wang et al., 2020; Zhang et al., 2024), since we are focused on the zero-shot setting, as well as sparsity/LSH-based approaches (Kitaev et al., 2020; Daras et al., 2020; Chen et al., 2021; Han et al., 2024), since these do not admit efficient implementations on current GPUs. We also benchmark our fast implementation of MonarchAttention, comparing with FlashAttention-2 (Dao, 2023).
@@
We convert all 12 attention layers, each having 12 heads with sequence length $N=197$ and head dimension $d=64$, of the 87M parameter ViT-B (Dosovitskiy et al., 2021) that has been pre-trained on ImageNet-21K (Deng et al., 2009) and fine-tuned on ImageNet-1K (Russakovsky et al., 2015) for image classification. To evaluate the performance at different FLOP counts, we vary the number of steps $T$ for MonarchAttention, and vary the rank for Performer and Nyströmformer; see Section D.2 for more details on the set-up. The results are shown in the left panel of Figure 2. MonarchAttention achieves significant improvement over other baselines – compared to the original softmax attention, MonarchAttention loses only 5% accuracy to reduce attention FLOPs by 80%, or matches the performance to reduce attention FLOPs by 50%.
@@
We convert the initial 4 and final 4 layers of the 12 attention layers, each having 12 heads with sequence length $N=384$ and head dimension $d=64$, of the 125M parameter RoBERTa-B (Liu et al., 2019) that has been pre-trained on a large English corpus and fine-tuned on SQuAD1.1 (Rajpurkar et al., 2016) for question answering. As before, to evaluate the performance at different FLOP counts, we vary the block size $b$ for MonarchAttention, and vary the rank for Performer and Nyströmformer; see Section D.3 for more details on the set-up. The results are shown in the right panel of Figure 2. Once again, MonarchAttention achieves significant improvement over other baselines – compared to the original softmax attention, MonarchAttention loses only 10 points in F1 score to reduce attention FLOPs by 60%, or matches the performance to reduce attention FLOPs by 35%.
@@
We convert all 6 attention layers, each having 12 heads with head dimension $d=64$, in the encoder of the 139M parameter BART-B (Lewis et al., 2020) that has been pre-trained on a large English corpus and fine-tuned on BookSum-chapters (Kryściński et al., 2022) for summarization. We only convert the encoder model and leave the decoder intact. To evaluate the benefits of sub-quadratic attention for processing longer sequences, we truncate the text to be summarized to various sequence lengths $N$ for each method; see Section D.4 for more details on the set-up. The results are shown in Figure 3. We see that MonarchAttention achieves a strictly better ROUGE score (Lin, 2004) vs. FLOPs tradeoff than even softmax attention, due to accurate and efficient processing of longer sequences. In particular, the $N=8192$ MonarchAttention model improves on the $N=2048$ softmax attention model by $0.75$ on ROUGE-1 and $0.5$ on ROUGE-L with slightly fewer FLOPs, while the $N=8192$ Nyströmformer model with similar FLOPs does strictly worse than softmax.
@@
We convert a subset of the 28 attention layers, each having 16 heads with sequence length $N=256$ and head dimension $d=72$, of the 675M parameter DiT-XL (Peebles and Xie, 2023) that has been trained on ImageNet (Deng et al., 2009). We consider replacing either all layers, the first $14$ layers, or the last $14$ layers; see Section D.5 for more details on the set-up. Examples of generated images with each method for replacing the first 14 layers are shown in Figure 5, where MonarchAttention produces clear images resembling those of softmax attention, while the Nyströmformer ones are extremely noisy. We also quantitatively evaluate the (s)FID scores (Heusel et al., 2017) of MonarchAttention compared with Nyströmformer, using images generated with the original softmax attention model as reference – the results are reported in Table 1. MonarchAttention noticeably outperforms Nyströmformer with similar FLOP count. In particular, using MonarchAttention in the first half of the DiT layers results in extremely small FID and sFID from the softmax attention model’s images, while reducing FLOPs by nearly $30\%$.

### v2

@@ Abstract
Transformers have achieved state-of-the-art performance across various tasks, but suffer from a notable quadratic complexity in sequence length due to the attention mechanism. In this work, we propose MonarchAttention – a novel approach to sub-quadratic attention approximation via Monarch matrices, an expressive class of structured matrices. Based on the variational form of softmax, we describe an efficient optimization-based algorithm to compute an approximate projection of softmax attention onto the class of Monarch matrices with $\Theta(N\sqrt{N}d)$ computational complexity and $\Theta(Nd)$ memory/IO complexity. Unlike previous approaches, MonarchAttention is both (1) transferable, yielding minimal performance loss with no additional training, even when replacing every attention layer of the Transformer, and (2) hardware-efficient, utilizing the highest-throughput tensor core units on modern GPUs. With optimized kernels, MonarchAttention achieves substantial speed-ups in wall-time over FlashAttention-2: $1.4\times$ for shorter sequences $(N=256)$, $4.5\times$ for medium-length sequences $(N=4K)$, and $8.2\times$ for longer sequences $(N=16K)$. We demonstrate the quality of MonarchAttention on diverse tasks and architectures in vision and language problems, showing that it flexibly and accurately approximates softmax attention in a variety of contexts. Our code is available at https://github.com/cjyaras/monarch-attention.
@@ 1 Introduction
Over the past decade, Transformers (Vaswani et al., 2017) have become the dominant architecture for generating and processing various data modalities, such as text (Brown et al., 2020), images (Dosovitskiy et al., 2021), and speech (Radford et al., 2023). Central to the Transformer’s success is attention, the mechanism through which complex interactions within sequential data are captured through weighted combinations of embeddings at every position in the sequence. Famously, the attention mechanism has a quadratic-time complexity $\Theta(N^{2}d)$ in the length of the sequence $N$, where $d$ is the head dimension, which is a key bottleneck for both training and inference, particularly in long sequence problems. To address this, numerous works have proposed sub-quadratic substitutes for attention. Yet, such approaches are either (1) not transferable, requiring training from scratch or fine-tuning of existing models, or (2) do not yield speed-ups in practice (except on extremely long sequences) due to a gap between theoretical complexity and practical considerations for modern GPUs, especially compared to highly optimized implementations (Dao et al., 2022b).
In this work, we propose MonarchAttention: a novel sub-quadratic attention substitute based on approximating the attention matrix via Monarch matrices (Dao et al., 2022a), a class of expressive structured matrices. At first glance, this is computationally infeasible – for sequence length $N$, computing an exact projection onto the set of Monarch matrices has a super-quadratic $O(N^{2}\sqrt{N})$-time complexity, not to mention that we need to form the entire $N\times N$ attention matrix. Instead, we reframe the computation of the attention matrix as an optimization problem in terms of the variational form of softmax, and exploit low-dimensional structure in the variational objective when constrained to the set of Monarch matrices – this yields a sub-quadratic $\Theta(N\sqrt{N}d)$-time approximation, where $d$ is the head dimension. This approach combines two optimization-based improvements for its success: the variational form of the softmax nonlinearity with a structured factorization approach analogous to those for low-rank approximation of a matrix (Chi et al., 2019), where rather than computing a full SVD and truncating to the desired rank, one can more efficiently minimize a Frobenius norm objective constrained to the set of low-rank matrices. We briefly review prior work on structured matrices, including Monarch matrices, as well as existing approaches to efficient attention.
@@
We note that there are significant drawbacks to the approaches described above. Pure low-rank methods are often fast and hardware-friendly, but are typically not suitable as drop-in replacements for attention in pre-trained Transformers due to the prevalence of “strongly diagonal”, high-rank attention matrices where attention weights are concentrated locally in a sequence. Making up for this with a fixed sparsity pattern does not allow for data-dependent support of the attention matrix, necessary for zero-shot conversion. Finally, sparsity/LSH-based approaches that do not have a fixed sparsity pattern improve on accuracy over low-rank approximations but suffer from significant overhead due to GPU incompatibility. MonarchAttention, on the other hand, achieves the best of both worlds: it is fast and hardware-friendly due to utilization of tensor cores for batched matmuls, while computing highly accurate approximations to the extent that it can directly replace softmax attention with no additional training.
@@ 2 Preliminaries
We use $[N]$ to denote the index set $\{1,2,\dots,N\}$. We use $\Delta^{N}$ to denote the (N1)(N-1) dimensional unit simplex, given by $\Delta^{N}=\{\bm{a}\in\mathbb{R}^{N}:\bm{a}\succeq 0,\langle\bm{1}_{N},\bm{a}\rangle=1\}$. We denote the $m\times m$ identity matrix by $\bm{I}_{m}$. We use the notation $\bm{A}_{ijk}$ to denote an element of a 3-way tensor, and $\bm{A}_{i,:,k}$ to denote a slice. We use $\delta_{kl}$ to denote the Kronecker delta that is $1$ if $k=l$ and otherwise 0.
@@
The softmax function $\mathbb{R}^{N}\rightarrow\Delta^{N}$ maps $N$ real numbers to the (N1)(N-1)-dimensional unit simplex, and is defined as
$\operatorname{softmax}(\bm{z})]_{i}:=\frac{\exp(\bm{z}_{i})}{\sum_{j}\exp(\bm{z}_{j})},\;\forall i\in[N].$ (1)
@@ 3 MonarchAttention
where $\Delta^{N\times N}$ denotes a matrix whose rows lie on $\Delta^{N}$, and $H(\bm{A})=-\sum_{i,j}\bm{A}_{ij}\log\bm{A}_{ij}$. Here we demonstrate how to handle this objective function when $\bm{A}$ has Monarch structure, and defer discussion of the simplex constraint to the next paragraph. For a dense matrix $\bm{A}$, computing $f(\bm{A};\bm{Q},\bm{K})$ requires $\Theta(N^{2}d)$ operations, which is the same as computing $\sigma(\bm{Q}\bm{K}^{\top})$ directly. However, when $\bm{A}$ is a Monarch matrix $\bm{M}=\bm{P}^{\top}\bm{B}$, we have
$f(\bm{P}^{\top}\bm{B};\bm{Q},\bm{K})$ $=\langle\bm{P}^{\top}\bm{B},\bm{Q}\bm{K}^{\top}\rangle+H(\bm{P}^{\top}\bm{B})$
$=\langle\bm{B},\widetilde{\bm{Q}}\bm{K}^{\top}\rangle+H(\bm{B})=\sum_{j,k}f(\bm{B}_{jk};\widetilde{\bm{Q}}_{j},\bm{K}_{k}),$
where $\widetilde{\bm{Q}}=\bm{P}\bm{Q}$, and $\widetilde{\bm{Q}}_{j}\in\mathbb{R}^{m\times d},\bm{K}_{k}\in\mathbb{R}^{b\times d}$ are the $j$-th and $k$-th block of rows of $\widetilde{\bm{Q}},\bm{K}$ respectively. Then, for each $j\in[b],k\in[m]$ we evaluate $f$ on the rank-one matrix $\bm{B}_{jk}=\bm{L}_{jk}\bm{R}_{kj}^{\top}$:
@@
$=\langle\widetilde{\bm{Q}}_{j}^{\top}\bm{L}_{jk},\bm{K}_{k}^{\top}\bm{R}_{kj}\rangle+\left(\bm{1}^{\top}\bm{R}_{kj}\right)\cdot H(\bm{L}_{jk})+\left(\bm{1}^{\top}\bm{L}_{jk}\right)\cdot H(\bm{R}_{kj}).$
@@
i.e., slices of $\bm{L},\bm{R}$ individually lying on the unit simplex is sufficient to enforce the constraint on $\bm{M}$. This is easily seen from (5) – obviously if $\bm{L}_{jkl},\bm{R}_{kji}\geq 0$, then $\bm{M}_{lk}]_{ji}\geq 0$. Moreover, this also enforces the sum-to-one constraint, as rows of $\bm{M}$ sum as
@@
We now present the updates for $\bm{L},\bm{R}$. Initializing $\bm{L}_{jkl}^{(0)}=\delta_{kl}$ as identity, we have
@@
where $\overline{\bm{Q}}_{jl},\overline{\bm{K}}_{ki}\in\mathbb{R}^{d}$ are the $(b\cdot(l-1)+j)$-th and $(b\cdot(k-1)+i)$-th rows of $\bm{Q}$ and $\bm{K}$ respectively. The full derivation is provided in Section C.1. After $T$ steps, we obtain the final Monarch approximation $\bm{M}^{(T)}\approx\sigma(\bm{Q}\bm{K}^{\top})$ with factors $\bm{L}^{(T)}$ and $\bm{R}^{(T)}$, from which we output $\bm{M}^{(T)}\bm{V}$ using (4). A naïve implementation of the full algorithm is provided in Section C.2. We discuss in Section C.3 how padding can be incorporated into MonarchAttention for when $N$ is not divisible by $b$. We empirically demonstrate the fast convergence of MonarchAttention in Section E.1.
@@
def al_cl_kernel(aR, cR, Kb): # Computes aL, cL from aR, cR
R = softmax(bmm(aR, Kb.transpose(1, 2)) / cR[:, :, None], dim=2)
cL = sum(R * log(R), dim=2).transpose(0, 1)
aL = bmm(R, Kb).transpose(0, 1)
return aL, cL
[…]
L = softmax(bmm(aL, Qb.transpose(1, 2)) - cL[:, :, None], dim=1)
cR = sum(L, dim=2).transpose(0, 1)
aR = bmm(L, Qb).transpose(0, 1)
return aR, cR
[…]
R = softmax(bmm(aR, Kb.transpose(1, 2)) / cR[:, :, None], dim=2)
cL = sum(R * log(R), dim=2).transpose(0, 1)
aL = bmm(R, Kb).transpose(0, 1)
y = bmm(R, Vb).transpose(0, 1)
return aL, y, cL
[…]
L = softmax(bmm(Qb, aL.transpose(1, 2)) - cL[:, None, :], dim=2)
z = bmm(L, y).transpose(0, 1)
return z
[…]
Qb = Q.reshape(m, b, d).transpose(0, 1)
Kb = K.reshape(m, b, d)
Vb = V.reshape(m, b, d)
aR = Q.reshape(m, b, d)
cR = ones(m, b)
[…]
aL, cL = al_cl_kernel(aR, cR, Kb)
aR, cR = ar_cr_kernel(aL, cL, Qb)
[…]
z = z_kernel(aL, y, cL, Qb)
o = z.reshape(N, d)
return o
@@
which can be seen as a batched attention computation, meaning we can implement a FlashAttention-like kernel to reduce IO between HBM and SRAM. However, several aspects of this computation make it particularly IO-efficient. Besides the fact that $\overline{\bm{K}}$ acts as both the $\bm{K}$ and $\bm{V}$ matrices in (3), the effective sequence length is $\sqrt{N}$. This eliminates the need for tiling along the sequence length, except for very long sequences having $\Theta(\sqrt{N}d)>S$, where $S$ is the size of on-chip SRAM. This means that we have an optimal IO complexity of $\Theta(Nd)$ for a single call, as opposed to the worst-case $O(N^{2}d^{2}/S)$ complexity of FlashAttention. The computation of $\bm{\alpha}_{R}^{(t)}$ from $\bm{\alpha}_{L}^{(t)}$, as well as the Monarch matmul, can be written in a similar fashion. Based on this, MonarchAttention not only achieves significant speed-up over FlashAttention for longer sequences, but also for shorter ones. Python-like code for MonarchAttention is given in Figure 4.
@@ 4 Experiments
In this section, we evaluate the zero-shot performance (no additional training) of MonarchAttention for converting pre-trained/fine-tuned Transformer attention layers to sub-quadratic attention in four different model/task settings. We compare with previous low-rank attention methods (Katharopoulos et al., 2020; Choromanski et al., 2021; Xiong et al., 2021; Qin et al., 2022); see Section D.1 for more details on the baselines. We specifically exclude low-rank methods with learnable components (Wang et al., 2020; Zhang et al., 2024), since we are focused on the zero-shot setting, as well as sparsity/LSH-based approaches (Kitaev et al., 2020; Daras et al., 2020; Chen et al., 2021; Han et al., 2024), since these do not admit efficient implementations on current GPUs. We also benchmark our fast implementation of MonarchAttention, comparing with FlashAttention-2 (Dao, 2024).
@@
We convert all 12 attention layers, each having 12 heads with sequence length $N=197$ and head dimension $d=64$, of the 87M parameter ViT-B (Dosovitskiy et al., 2021) that has been pre-trained on ImageNet-21K (Deng et al., 2009) and fine-tuned on ImageNet-1K (Russakovsky et al., 2015) for image classification. To evaluate the performance at different FLOP counts, we vary the number of steps $T\in\{1,2,3\}$ for MonarchAttention, and vary the rank for Performer and Nyströmformer; see Section D.2 for more details on the set-up. The results are shown in the left panel of Figure 2. MonarchAttention achieves significant improvement over other baselines – compared to the original softmax attention, MonarchAttention loses only 5% accuracy to reduce attention FLOPs by 80%, or matches the performance to reduce attention FLOPs by 50%.
@@
We convert the initial 4 and final 4 layers of the 12 attention layers, each having 12 heads with sequence length $N=384$ and head dimension $d=64$, of the 125M parameter RoBERTa-B (Liu et al., 2019) that has been pre-trained on a large English corpus and fine-tuned on SQuAD1.1 (Rajpurkar et al., 2016) for question answering. To evaluate the performance at different FLOP counts, we fix $T=1$ and vary the block size $b$ for MonarchAttention, and vary the rank for Performer and Nyströmformer; see Section D.3 for more details on the set-up. The results are shown in the right panel of Figure 2. Once again, MonarchAttention achieves significant improvement over other baselines – compared to the original softmax attention, MonarchAttention loses only 10 points in F1 score to reduce attention FLOPs by 60%, or matches the performance to reduce attention FLOPs by 35%. We additionally provide a per-layer ablation study in Section E.2 to understand the impact of converting different layers in this setting.
@@
We convert all 6 attention layers, each having 12 heads with head dimension $d=64$, in the encoder of the 139M parameter BART-B (Lewis et al., 2020) that has been pre-trained on a large English corpus and fine-tuned on BookSum-chapters (Kryściński et al., 2022) for summarization. We only convert the encoder model and leave the decoder intact. To evaluate the benefits of sub-quadratic attention for processing longer sequences, we truncate the text to be summarized to various sequence lengths $N$ for each method, with $T\in\{2,3\}$ for MonarchAttention depending on the sequence length; see Section D.4 for more details on the set-up. The results are shown in Figure 3. We see that MonarchAttention achieves a strictly better ROUGE score (Lin, 2004) vs. FLOPs tradeoff than even softmax attention, due to accurate and efficient processing of longer sequences. In particular, the $N=8192$ MonarchAttention model improves on the $N=2048$ softmax attention model by $0.75$ on ROUGE-1 and $0.5$ on ROUGE-L with slightly fewer FLOPs, while the $N=8192$ Nyströmformer model with similar FLOPs does strictly worse than softmax. Although we are mainly focused on the zero-shot setting, we also investigate the post-swap fine-tuning performance and stability of MonarchAttention on this task in Section E.3.
@@
We convert a subset of the 28 attention layers, each having 16 heads with sequence length $N=256$ and head dimension $d=72$, of the 675M parameter DiT-XL (Peebles and Xie, 2023) that has been trained on ImageNet (Deng et al., 2009). We consider replacing either all layers, the first $14$ layers, or the last $14$ layers, and fix $T=3$ for MonarchAttention; see Section D.5 for more details on the set-up. Examples of generated images with each method for replacing the first 14 layers are shown in Figure 5, where MonarchAttention produces clear images resembling those of softmax attention, while the Nyströmformer ones are extremely noisy. We also quantitatively evaluate the (s)FID scores (Heusel et al., 2017) of MonarchAttention compared with Nyströmformer, using images generated with the original softmax attention model as reference – the results are reported in Table 1. MonarchAttention noticeably outperforms Nyströmformer with similar FLOP count. In particular, using MonarchAttention in the first half of the DiT layers results in extremely small FID and sFID from the softmax attention model’s images, while reducing FLOPs by nearly $30\%$. We provide a more fine-grained version of Table 1 for MonarchAttention in Section E.4 to investigate the impact of replacing different quarters of the DiT model.
Node Classification with Graph Transformer.
Beyond image and language tasks, we also evaluate the performance of MonarchAttention for converting attention in Graph Transformers on a graph node classification task in Section E.5.